A box with no top is to be constructed from a piece of cardboard whose length measures 12 inches more than its width. The box is formed by cutting squares that measure 5 inch on each side from the four corners and then folding up the sides.
If the volume of the box 65 in ^3 what are the dimensions of the piece of cardboard?
The Height?
The width?

Answer by TimothyLamb(4379)

---
v = 65 = Lwh
h = 5
65 = 5Lw
---
Lw = 13
L = w + 12
---
Lw = 13
(w + 12)w = 13
ww + 12w - 13 = 0
---
the above quadratic equation is in standard form, with a=1, b=12, and c=-13
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 12 -13
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
w = 1
w = -13
---
negative width doesn't make sense so use the positive root:
---
dimensions of the bottom of the box
w = 1
L = w + 12 = 13
---
dimensions of the cardboard:
w= 1 + 10 = 11 inch
L= 13 + 10 = 23 inch
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php